a spring has a force of 80 n/m, it vibrates with a frequency of 7 Hz, Calculate the, a) period b)angular frequency; c) mass of the spring.


1
Expert's answer
2021-02-15T17:03:27-0500

a)

"T=\frac{1}{\nu},~~T=\frac 17\approx 0.14 ~\text{s},"

b)

"\omega=2\pi\nu,~~\omega=2\cdot3.14\cdot7\approx44~\frac{\text{rad}}{\text{s}},"

c)

"\omega=\sqrt{\frac km},~~m=\frac{k}{\omega^2},~~m=\frac{80}{44^2}\approx41~\text{g}."